Given the Production Function Q = 72X + 15X2 - X3, where Q =Output and X=Input

a) What is the Marginal Product (MP) when X = 8?

b) What is the Average Product (AP) when X = 6?

c) At what value of X will Q be at its maximum?

d) At what value of X will Diminishing Returns set in?


1
Expert's answer
2021-05-26T11:50:38-0400

Given "72X + 15X^2 - X^3"


a) "MP=\frac {dQ} {dX} =72+30X-3X^2"


When X=8;


"72+30(8)-3(8)^2"


=72+240−192=120


MP=120


b) "AP=\frac {Q} {X} =72+15X-X^2"


When X=6;


AP=72+15(6)−(6×6)=126


AP=126


c) Q is maximum when "\frac {dQ} {dX} =0"


"72+30X-3X^2=0"


"3X^2-30X-72=0"


"X^2-10X-24=0"


"X^2-12X+2X-24=0"


"X(X\u221212)+2(X\u221212)=0"


X=12or (X=−2 which is inadmissible)


d) Diminishing returns means, MP is decreasing i.e

"\frac {dMP} {dX} \lt 0"


"\frac {dMP} {dX} =0"


30-9X=0


9X=30


X=3.33


Thus diminishing returns sets is when X>3.33